OpenStudy (anonymous):

WHAT IS 1+1 HELP PLEASE I FORGOT HOW TO ADD

OpenStudy (anonymous):

ur funny

OpenStudy (tanya123):

1+1 = 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@Chocolate1031 SO DO U KNOW THE ANSWER

OpenStudy (anonymous):

ta I think it's window

OpenStudy (anonymous):

no it's 5

OpenStudy (anonymous):

ARE U GUYS SURE ITS 2 I GOT 9

OpenStudy (anonymous):

oops it's actually 50

OpenStudy (anonymous):

hahaha

OpenStudy (tanya123):

Lmao

OpenStudy (anonymous):

@cram can u explain

OpenStudy (tanya123):

He can explain you @phi

OpenStudy (anonymous):

@tanya123 can u explain how you got that answer please

OpenStudy (anonymous):

1+1=2 okay one diamond plus another equals two diamonds

OpenStudy (anonymous):

1+1=2

OpenStudy (tanya123):

You can use a calculator

OpenStudy (anonymous):

@cram i dont have any diamonds but i have two HOT DOGS will that work

OpenStudy (anonymous):

u guys are all wrong, because 1 + 1 = two.. like duh

OpenStudy (anonymous):

lmao so hard please stop

OpenStudy (anonymous):

wat is calculator

OpenStudy (anonymous):

a calculator is

OpenStudy (anonymous):

is that a type of hot dogs because hot dogs are the only thing i have in life :(

OpenStudy (tanya123):

something used for making mathematical calculations, in particular a small electronic device with a keyboard and a visual display.

OpenStudy (anonymous):

This is really hard @Kidz_Wavy , why don't I call my baby brother who is FYI 1 yrs old, and tell him to answer

OpenStudy (anonymous):

jking but hmu on kik and Instagram @kidz_wavy

OpenStudy (anonymous):

@cram

OpenStudy (anonymous):

no no wait, shouldn't 1+1 be 11, 1, or 0

OpenStudy (anonymous):

lmao my stomach hurts now

OpenStudy (anonymous):

btw my brother said it should be his toy

OpenStudy (anonymous):

@Chocolate1031 i dont think my hot dogs can call people but i will try can u give me your baby bros #

OpenStudy (anonymous):

ya its 911

OpenStudy (anonymous):

lmao chocalte this is freaking funny

OpenStudy (anonymous):

oops wait that's my number, his number is 1-800-LADYBUGSARECOOL

OpenStudy (anonymous):

i called and thay said they with send a dispatch team to my house to help me thanks @Chocolate1031

OpenStudy (anonymous):

then y are u sending a dispatch team to my house ?

OpenStudy (anonymous):

umm what

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

LMAO

OpenStudy (anonymous):

@cram give me ur kik plz mine is KIDZ_WAVY

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

don't have one sorry

OpenStudy (anonymous):

@cram k than be racist

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

not evan a instagram or facebook

OpenStudy (anonymous):

i am not racist i like all races

OpenStudy (anonymous):

is it because im blue

OpenStudy (anonymous):

im orange

OpenStudy (anonymous):

MY MOM is orange 2

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

thats amazing